# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 3.3 Bounded Real Lemma in Contraction Theory
Setting $\kappa=\delta q$ and viewing $\varsigma$ as external disturbance, this system can be interpreted as the differential closed-loop dynamics defined earlier in (3.20). The bounded real lemma states that this system is $\mathcal{L}_{2}$ gain stable with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$, i.e., $\|(\upsilon)_{\tau}\|_{\mathcal{L}_{2}}\leq\gamma\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}+const.$, or equivalently, the $\mathcal{H}_{\infty}$ norm of the transfer function of (3.37) is less than or equal to $\gamma$, if the following LMI for $\mathcal{P}\succ 0$ holds (see [Başar and Bernhard, 2008, p. 369]):
@@
Theorem 3.3 introduces the bounded real lemma in the context of contraction theory.
@@
Let $\mathcal{A}=A-BR^{-1}B^{\top}M$, $\mathcal{B}=M^{-1}$, $C=\sqrt{2\alpha}\Theta$, and $\mathcal{D}=0$ in (3.37), where $M=\Theta^{\top}\Theta$, and the other variables are as defined in Theorem 3.1. Then (3.38) with $\mathcal{P}=M$ and $\gamma=1/\sqrt{\beta}$ is equivalent to (3.25), and thus (3.25) implies $\mathcal{L}_{2}$ gain stability of (3.37) with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$.
@@
resulting in $\dot{V}+\|\upsilon\|^{2}-\gamma^{2}\|\varsigma\|^{2}\preceq 0$ for $V=\kappa^{\top}\mathcal{P}\kappa=\delta q^{\top}M\delta q$. This implies $\mathcal{L}_{2}$ gain stability with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma=1/\sqrt{\beta}$ [Khalil, 2002, p. 209]. $\qed$
3.4 KYP Lemma in Contraction Theory
Analogously to Theorem 3.3, the LMI (3.25) of Theorem 3.1 can be understood using the Kalman-Yakubovich-Popov (KYP) lemma [Dullerud and Paganini, 2000, p. 218]. Consider the quadratic Lyapunov function $V=\kappa^{\top}\mathcal{Q}\kappa$ with $\mathcal{Q}\succ 0$, satisfying the following output strict passivity (dissipativity) condition:
@@ 3.4 KYP Lemma in Contraction Theory
This implies that we have $\|(\upsilon)_{\tau}\|_{\mathcal{L}_{2}}\leq\gamma\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}+\sqrt{\gamma V(0)}$ by the comparison lemma [Khalil, 2002, pp. 102-103, pp. 350-353], leading to $\mathcal{L}_{2}$ gain stability with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$ [Dullerud and Paganini, 2000, p. 218]. The condition (3.40) can be expressed equivalently as an LMI form as follows:
@@
If (3.42) holds, the LMI (3.38) for the bounded real lemma holds with $\mathcal{P}=\gamma\mathcal{Q}$, i.e., the system (3.37) is $\mathcal{L}_{2}$ gain stable with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$. Thus, for systems with $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$ defined in Theorem 3.3, the condition (3.42) guarantees the contraction condition (3.25) of Theorem 3.1.
@@
Therefore, a necessary condition of (3.42) reduces to (3.38) if $\mathcal{Q}=\mathcal{P}/\gamma$. The rest follows from Theorem $3.$ $\qed$
@@
Theorems 3.3 and 3.4 imply the following statements.
@@
  • (3.40) is equivalent to (3.42), and to the output strict passivity condition with dissipation $1/\gamma$ [Khalil, 2002, p. 231].
@@
4 Convex Optimality in Robust Nonlinear Control and Estimation via Contraction Theory
Theorem 3.1 indicates that the problem of finding contraction metrics for general nonlinear systems could be formulated as a convex feasibility problem. This section thus delineates one approach, called the method of ConVex optimization-based Steady-state Tracking Error Minimization (CV-STEM) Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019], to optimally design $M$ of Theorem 3.1 that defines a contraction metric and minimizes an upper bound of the steady-state tracking error in Theorem 2.4 or in Theorem 2.5 via convex optimization. In particular, we present an overview of the SDC- and CCM-based CV-STEM frameworks for provably stable and optimal feedback control and state estimation of nonlinear systems, perturbed by deterministic and stochastic disturbances. It is worth noting that the steady-state bound is expressed as a function of the condition number $\chi=\overline{m}/\underline{m}$ as to be seen in (4.1) and (4.19), which renders the CV-STEM applicable and effective even to learning-based and data-driven control frameworks as shall be seen in Sec. 5.
4.1 CV-STEM Control
As a result of Theorems 2.42.5, and 3.1, the control law $u=u_{d}-K(x-x_{d})$ of (3.15) gives a convex steady-state upper bound of the Euclidean distance between the system trajectories, which can be used as an objective function for the CV-STEM control framework in Theorem 4.2 Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019].
@@ 4.1 CV-STEM Control
If one of the matrix inequalities of Theorem 3.1 holds, we have the following bound for $\chi={\overline{m}}/{\underline{m}}$:
@@
where $c_{0}(\alpha,\alpha_{G})=\bar{d}_{c}/\alpha$ for deterministic systems and $c_{0}(\alpha,\alpha_{G})=\bar{g}_{c}\sqrt{(2\alpha_{G}^{-1}+1)/(2\alpha)}$ for stochastic systems, with the variables given in Theorem 3.1.
@@
Taking $\lim_{t\to\infty}$ in the exponential tracking error bounds (3.27) and (3.28) of Theorem 3.1 gives the first inequality of (4.1). The second inequality follows from the relation $1\leq\sqrt{\overline{m}/\underline{m}}\leq\overline{m}/\underline{m}=\chi$ due to $\underline{m}\leq\overline{m}$. $\qed$
The convex optimization problem of minimizing (4.1), subject to the contraction constraint of Theorem 3.1, is given in Theorem 4.2, thereby introducing the CV-STEM control Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019] for designing an optimal contraction metric and contracting control policy as depicted in Fig. 2.
@@
Suppose that $\alpha$, $\alpha_{G}$, $\bar{d}_{c}$, and $\bar{g}_{c}$ in (4.1) and the Lipschitz constant $L_{m}$ of $\partial M/\partial x_{i}$ in Theorem 3.1 are given. If the pair $(A,B)$ is uniformly controllable, the non-convex optimization problem of minimizing the upper bound (4.1) is equivalent to the following convex optimization problem, with the convex contraction constraint (3.24) or (3.25) of Theorem 3.1 and $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
where $W=M(x,x_{d},u_{d},t)^{-1}$ for Theorem 3.1 or $W=M(x,t)^{-1}$ for Theorem 3.2, $\bar{W}=\nu W$, $\nu=\overline{m}$, $\chi=\overline{m}/\underline{m}$, $c_{0}$ is as defined in (4.1) of Theorem 4.1, $c_{1},c_{2}\in\mathbb{R}_{\geq 0}$, and $P$ is a performance-based convex cost function.
@@
As proven in Theorems 3.1 and 4.1, $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), the contraction constraint (3.22), and the objective (4.1) reduce to (4.2) with $c_{1}=0$ and $c_{2}=0$. Since the resultant constraints are convex and the objective is affine in terms of the decision variables $\nu$, $\chi$, and $\bar{W}$, the problem (4.2) is indeed convex. Since we have $\sup_{x,t}\|M\|\leq\overline{m}=\nu$, $c_{1}$ can be used as a penalty to optimally adjust the induced $2$-norm of the control gain (see Example 4.1). The SDC coefficients $\varrho$ can also be utilized as decision variables for controllability due to Proposition 1 of Tsukamoto and Chung [2021]. $\qed$
@@
The weights $c_{0}$ and $c_{1}$ of the CV-STEM control of Theorem 4.2 establish an analogous trade-off to the case of the LQR with the cost weight matrices of $Q$ for state and $R$ for control, since the upper bound of the steady-state tracking error (4.1) is proportional to $\chi$, and an upper bound of $\|K\|$ for the control gain $K$ in (3.15), $u=u_{d}-K(x-x_{d})$, is proportional to $\nu$. In particular Tsukamoto and Chung [2021], Tsukamoto et al. [2021],
  • if $c_{1}$ is much greater than $c_{0}$ in (4.2) of Theorem 4.2, we get smaller control effort but with a large steady-state tracking error, and
  • if $c_{1}$ is much smaller than $c_{0}$ in (4.2) of Theorem 4.2, we get a smaller steady-state state tracking error but with larger control effort.
This is also because the solution $P\succ 0$ of the LQR Riccati equation [Rugh, 1996, p. 89], $\dot{P}=PA+A^{\top}P-PBR^{-1}B^{\top}P+Q$, can be viewed as a positive definite matrix $M$ that defines a contraction metric as discussed in Example 2.3.
@@
4.2 CV-STEM Estimation
We could also design an optimal state estimator analogously to the CV-STEM control of Theorem 4.2, due to the differential nature of contraction theory that enables LTV systems-type approaches to stability analysis. In particular, we exploit the estimation and control duality in differential dynamics similar to that of the Kalman filter and LQR in LTV systems.
@@ 4.2 CV-STEM Estimation
where $\bar{C}(\varrho_{c},\hat{x},t)=C(\varrho_{c},\hat{x},\bar{x},t)$ for a fixed trajectory $\bar{x}$ (e.g., $\bar{x}=0$, see Theorem 3.2), $R(\hat{x},t)\succ 0$ is a weight matrix on the measurement $y$, and $M(\hat{x},t)\succ 0$ is a positive definite matrix (which satisfies the matrix inequality constraint for a contraction metric, to be given in (4.13) of Theorem 4.5). Note that we could use other forms of estimation laws such as the EKF Bonnabel and Slotine [2015], Tsukamoto and Chung [2021], Brown and Hwang [1997], analytical SLAM Tan et al. [2017], or SDC with respect to a fixed point Dani et al. [2015], Tsukamoto and Chung [2021, 2019], depending on the application of interest, which result in a similar stability analysis as in Theorem 3.2.
4.2.1 Nonlinear Stability Analysis of SDC-based State Estimation using Contraction Theory
@@
The similarity between (3.20) ($\partial_{\mu}\dot{q}=(A-BK)\partial_{\mu}q$) and (4.11) leads to the following theorem Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019]. Again, note that we could also use the SDC formulation with respect to a fixed point as delineated in Theorem 3.2 and as demonstrated in Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019].
@@
Suppose $\exists\bar{\rho},\bar{c}\in\mathbb{R}_{\geq 0}$ s.t. $\|R^{-1}(\hat{x},t)\|\leq\bar{\rho}$, $\|C(\varrho_{c},x,\hat{x},t)\|\leq\bar{c},~\forall x,\hat{x},t$. Suppose also that $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}I$ of (2.32) holds, or equivalently, $\mathrm{I}\preceq\bar{W}\preceq\chi I$ of (3.26) holds with $W=M(\hat{x},t)^{-1}$, $\bar{W}=\nu W$, $\nu=\overline{m}$, and $\chi=\overline{m}/\underline{m}$. As in Theorem 3.1, let $\beta$ be defined as $\beta=0$ for deterministic systems (4.3) and
@@
for stochastic systems (4.4), where $2\alpha_{e0}=L_{m}\bar{g}_{e0}^{2}(\alpha_{G}+{1}/{2})$, $2\alpha_{e1}=L_{m}\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}(\alpha_{G}+{1}/{2})$, $L_{m}$ is the Lipschitz constant of $\partial W/\partial x_{i}$, $\bar{g}_{e0}$ and $\bar{g}_{e1}$ are given in (4.4), and $\exists\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem 2.5.
@@
then Theorems 2.4 and 2.5 hold for the virtual systems (4.8) and (4.9), respectively, i.e., we have the following bounds for $\mathtt{e}=\hat{x}-x$ with $\nu=\overline{m}$ and $\chi=\overline{m}/\underline{m}$:
@@
where $V_{s\ell}=\int^{\hat{x}}_{x}\delta q^{\top}W\delta q$ and $V_{\ell}=\int^{\hat{x}}_{x}\|\Theta\delta q\|$ are given in Theorem 2.3 with $W=M^{-1}=\Theta^{\top}\Theta$ defining a contraction metric, the disturbance bounds $\bar{d}_{e0}$, $\bar{d}_{e1}$, $\bar{g}_{e0}$, and $\bar{g}_{e1}$ are given in (4.3) and (4.4), respectively, $C_{e0}=\bar{g}_{e0}^{2}({2}{\alpha_{G}}^{-1}+1)$, and $C_{e1}=\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}({2}{\alpha_{G}}^{-1}+1)$. Note that for stochastic systems, the probability that $\|\mathtt{e}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Theorem 3.1 indicates that (4.13) is equivalent to
@@
for $d$ in (4.8) and $G$ in (4.9), the bounds (4.14)–(4.16) follow from the proofs of Theorems 2.4 and 2.5 Tsukamoto and Chung [2021], Tsukamoto et al. [2021]. $\qed$
@@
Although (4.13) is not an LMI due to the nonlinear term $\nu\beta\mathrm{I}$ on its right-hand side for stochastic systems (4.4), it is a convex constraint as $\nu\beta=-\nu\alpha_{s}=-\nu\alpha_{e0}-\nu^{3}\alpha_{e1}$ is a concave function for $\nu\in\mathbb{R}_{>0}$ Boyd and Vandenberghe [2004], Tsukamoto et al. [2021].
4.2.2 CV-STEM Formulation for State Estimation
The estimator (4.7) gives a convex steady-state upper bound of the Euclidean distance between $x$ and $\hat{x}$ as in Theorem 4.1 Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019].

### v8

@@ 3.3. Bounded Real Lemma in Contraction Theory
Setting $\kappa=\delta q$ and viewing $\varsigma$ as external disturbance, this system can be interpreted as the differential closed-loop dynamics defined earlier in (3.20). The bounded real lemma states that this system is $\mathcal{L}_{2}$ gain stable with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$, i.e., $\|(\upsilon)_{\tau}\|_{\mathcal{L}_{2}}\leq\gamma\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}+const.$, or equivalently, the $\mathcal{H}_{\infty}$ norm of the transfer function of (3.37) is less than or equal to $\gamma$, if the following LMI for $\mathcal{P}\succ 0$ holds (see [brlbook, p. 369]):
@@
Theorem LABEL:THM:Thm:bounded_real_cvstem introduces the bounded real lemma in the context of contraction theory.
@@
Bounded Real Lemma in Contraction TheoryThm:bounded_real_cvstem Let $\mathcal{A}=A-BR^{-1}B^{\top}M$, $\mathcal{B}=M^{-1}$, $C=\sqrt{2\alpha}\Theta$, and $\mathcal{D}=0$ in (3.37), where $M=\Theta^{\top}\Theta$, and the other variables are as defined in Theorem LABEL:THM:Thm:CVSTEM:LMI. Then (3.38) with $\mathcal{P}=M$ and $\gamma=1/\sqrt{\beta}$ is equivalent to (3.25), and thus (3.25) implies $\mathcal{L}_{2}$ gain stability of (3.37) with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$.
@@
resulting in $\dot{V}+\|\upsilon\|^{2}-\gamma^{2}\|\varsigma\|^{2}\preceq 0$ for $V=\kappa^{\top}\mathcal{P}\kappa=\delta q^{\top}M\delta q$. This implies $\mathcal{L}_{2}$ gain stability with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma=1/\sqrt{\beta}$ [Khalil:1173048, p. 209]. $\qed$
3.4. KYP Lemma in Contraction Theory
Analogously to Theorem LABEL:THM:Thm:bounded_real_cvstem, the LMI (3.25) of Theorem LABEL:THM:Thm:CVSTEM:LMI can be understood using the Kalman-Yakubovich-Popov (KYP) lemma [kypbook, p. 218]. Consider the quadratic Lyapunov function $V=\kappa^{\top}\mathcal{Q}\kappa$ with $\mathcal{Q}\succ 0$, satisfying the following output strict passivity (dissipativity) condition:
@@ 3.4. KYP Lemma in Contraction Theory
This implies that we have $\|(\upsilon)_{\tau}\|_{\mathcal{L}_{2}}\leq\gamma\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}+\sqrt{\gamma V(0)}$ by the comparison lemma [Khalil:1173048, pp. 102-103, pp. 350-353], leading to $\mathcal{L}_{2}$ gain stability with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$ [kypbook, p. 218]. The condition (3.40) can be expressed equivalently as an LMI form as follows:
@@
KYP Lemma in Contraction TheoryThm:kyp_cvstem If (3.42) holds, the LMI (3.38) for the bounded real lemma holds with $\mathcal{P}=\gamma\mathcal{Q}$, i.e., the system (3.37) is $\mathcal{L}_{2}$ gain stable with its $\mathcal{L}_{2}$ gain less than or equal to $\gamma$. Thus, for systems with $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$ defined in Theorem LABEL:THM:Thm:bounded_real_cvstem, the condition (3.42) guarantees the contraction condition (3.25) of Theorem LABEL:THM:Thm:CVSTEM:LMI.
@@
Therefore, a necessary condition of (3.42) reduces to (3.38) if $\mathcal{Q}=\mathcal{P}/\gamma$. The rest follows from Theorem LABEL:THM:Thm:bounded_real_cvstem. $\qed$
@@
Contraction and Nonlinear Stability Theorems LABEL:THM:Thm:bounded_real_cvstem and LABEL:THM:Thm:kyp_cvstem imply the following statements.
@@
  • (3.40) is equivalent to (3.42), and to the output strict passivity condition with dissipation $1/\gamma$ [Khalil:1173048, p. 231].
@@
4. Convex Optimality in Contraction Theory
Theorem LABEL:THM:Thm:CVSTEM:LMI indicates that the problem of finding contraction metrics for general nonlinear systems could be formulated as a convex feasibility problem. This section thus delineates one approach, called the method of ConVex optimization-based Steady-state Tracking Error Minimization (CV-STEM) mypaperTAC, ncm, nscm, mypaper, to optimally design $M$ of Theorem LABEL:THM:Thm:CVSTEM:LMI that defines a contraction metric and minimizes an upper bound of the steady-state tracking error in Theorem LABEL:THM:Thm:Robust_contraction_original or in Theorem LABEL:THM:Thm:robuststochastic via convex optimization. In particular, we present an overview of the SDC- and CCM-based CV-STEM frameworks for provably stable and optimal feedback control and state estimation of nonlinear systems, perturbed by deterministic and stochastic disturbances. It is worth noting that the steady-state bound is expressed as a function of the condition number $\chi=\overline{m}/\underline{m}$ as to be seen in (4.1) and (4.19), which renders the CV-STEM applicable and effective even to learning-based and data-driven control frameworks as shall be seen in Sec. 5.
4.1. CV-STEM Control
As a result of Theorems LABEL:THM:Thm:Robust_contraction_originalLABEL:THM:Thm:robuststochastic, and LABEL:THM:Thm:CVSTEM:LMI, the control law $u=u_{d}-K(x-x_{d})$ of (3.15) gives a convex steady-state upper bound of the Euclidean distance between the system trajectories, which can be used as an objective function for the CV-STEM control framework in Theorem LABEL:THM:Thm:CV-STEM mypaperTAC, ncm, nscm, mypaper.
@@ 4.1. CV-STEM Control
Convexity of Steady-State Tracking Error BoundsThm:CVSTEMobjective If one of the matrix inequalities of Theorem LABEL:THM:Thm:CVSTEM:LMI holds, we have the following bound for $\chi={\overline{m}}/{\underline{m}}$:
@@
where $c_{0}(\alpha,\alpha_{G})=\bar{d}_{c}/\alpha$ for deterministic systems and $c_{0}(\alpha,\alpha_{G})=\bar{g}_{c}\sqrt{(2\alpha_{G}^{-1}+1)/(2\alpha)}$ for stochastic systems, with the variables given in Theorem LABEL:THM:Thm:CVSTEM:LMI.
@@
Taking $\lim_{t\to\infty}$ in the exponential tracking error bounds (3.27) and (3.28) of Theorem LABEL:THM:Thm:CVSTEM:LMI gives the first inequality of (4.1). The second inequality follows from the relation $1\leq\sqrt{\overline{m}/\underline{m}}\leq\overline{m}/\underline{m}=\chi$ due to $\underline{m}\leq\overline{m}$. $\qed$
The convex optimization problem of minimizing (4.1), subject to the contraction constraint of Theorem LABEL:THM:Thm:CVSTEM:LMI, is given in Theorem LABEL:THM:Thm:CV-STEM, thereby introducing the CV-STEM control mypaperTAC, ncm, nscm, mypaper for designing an optimal contraction metric and contracting control policy as depicted in Fig. 2.
@@
Robust Nonlinear Control via Convex OptimizationThm:CV-STEM Suppose that $\alpha$, $\alpha_{G}$, $\bar{d}_{c}$, and $\bar{g}_{c}$ in (4.1) and the Lipschitz constant $L_{m}$ of $\partial M/\partial x_{i}$ in Theorem LABEL:THM:Thm:CVSTEM:LMI are given. If the pair $(A,B)$ is uniformly controllable, the non-convex optimization problem of minimizing the upper bound (4.1) is equivalent to the following convex optimization problem, with the convex contraction constraint (3.24) or (3.25) of Theorem LABEL:THM:Thm:CVSTEM:LMI and $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
where $W=M(x,x_{d},u_{d},t)^{-1}$ for Theorem LABEL:THM:Thm:CVSTEM:LMI or $W=M(x,t)^{-1}$ for Theorem LABEL:THM:Thm:fixed_sdc, $\bar{W}=\nu W$, $\nu=\overline{m}$, $\chi=\overline{m}/\underline{m}$, $c_{0}$ is as defined in (4.1) of Theorem LABEL:THM:Thm:CVSTEMobjective, $c_{1},c_{2}\in\mathbb{R}_{\geq 0}$, and $P$ is a performance-based convex cost function.
@@
As proven in Theorems LABEL:THM:Thm:CVSTEM:LMI and LABEL:THM:Thm:CVSTEMobjective, $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), the contraction constraint (3.22), and the objective (4.1) reduce to (4.2) with $c_{1}=0$ and $c_{2}=0$. Since the resultant constraints are convex and the objective is affine in terms of the decision variables $\nu$, $\chi$, and $\bar{W}$, the problem (4.2) is indeed convex. Since we have $\sup_{x,t}\|M\|\leq\overline{m}=\nu$, $c_{1}$ can be used as a penalty to optimally adjust the induced $2$-norm of the control gain (see Example LABEL:EX:ex:lqrexample). The SDC coefficients $\varrho$ can also be utilized as decision variables for controllability due to Proposition 1 of mypaperTAC. $\qed$
@@
LQR Analogyex:lqrexample The weights $c_{0}$ and $c_{1}$ of the CV-STEM control of Theorem LABEL:THM:Thm:CV-STEM establish an analogous trade-off to the case of the LQR with the cost weight matrices of $Q$ for state and $R$ for control, since the upper bound of the steady-state tracking error (4.1) is proportional to $\chi$, and an upper bound of $\|K\|$ for the control gain $K$ in (3.15), $u=u_{d}-K(x-x_{d})$, is proportional to $\nu$. In particular ncm, nscm,
  • if $c_{1}$ is much greater than $c_{0}$ in (4.2) of Theorem LABEL:THM:Thm:CV-STEM, we get smaller control effort but with a large steady-state tracking error, and
  • if $c_{1}$ is much smaller than $c_{0}$ in (4.2) of Theorem LABEL:THM:Thm:CV-STEM, we get a smaller steady-state state tracking error but with larger control effort.
This is also because the solution $P\succ 0$ of the LQR Riccati equation [lssbook, p. 89], $\dot{P}=PA+A^{\top}P-PBR^{-1}B^{\top}P+Q$, can be viewed as a positive definite matrix $M$ that defines a contraction metric as discussed in Example LABEL:EX:ex:lyapunov_contraction.
@@
4.2. CV-STEM Estimation
We could also design an optimal state estimator analogously to the CV-STEM control of Theorem LABEL:THM:Thm:CV-STEM, due to the differential nature of contraction theory that enables LTV systems-type approaches to stability analysis. In particular, we exploit the estimation and control duality in differential dynamics similar to that of the Kalman filter and LQR in LTV systems.
@@ 4.2. CV-STEM Estimation
where $\bar{C}(\varrho_{c},\hat{x},t)=C(\varrho_{c},\hat{x},\bar{x},t)$ for a fixed trajectory $\bar{x}$ (e.g., $\bar{x}=0$, see Theorem LABEL:THM:Thm:fixed_sdc), $R(\hat{x},t)\succ 0$ is a weight matrix on the measurement $y$, and $M(\hat{x},t)\succ 0$ is a positive definite matrix (which satisfies the matrix inequality constraint for a contraction metric, to be given in (4.13) of Theorem LABEL:THM:Thm:CV-STEM-estimation). Note that we could use other forms of estimation laws such as the EKF 6849943, ncm, Brown:680442, analytical SLAM doi:10.1177/0278364917710541, or SDC with respect to a fixed point Ref:Stochastic, mypaperTAC, mypaper, depending on the application of interest, which result in a similar stability analysis as in Theorem LABEL:THM:Thm:fixed_sdc.
4.2.1. Nonlinear Stability Analysis of SDC-based State Estimation
@@
The similarity between (3.20) ($\partial_{\mu}\dot{q}=(A-BK)\partial_{\mu}q$) and (4.11) leads to the following theorem mypaperTAC, ncm, nscm, mypaper. Again, note that we could also use the SDC formulation with respect to a fixed point as delineated in Theorem LABEL:THM:Thm:fixed_sdc and as demonstrated in mypaperTAC, Ref:Stochastic, mypaper.
@@
Convexity of Contraction Conditions for EstimationThm:CVSTEM-LMI-est Suppose $\exists\bar{\rho},\bar{c}\in\mathbb{R}_{\geq 0}$ s.t. $\|R^{-1}(\hat{x},t)\|\leq\bar{\rho}$, $\|C(\varrho_{c},x,\hat{x},t)\|\leq\bar{c},~\forall x,\hat{x},t$. Suppose also that $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}I$ of (2.32) holds, or equivalently, $\mathrm{I}\preceq\bar{W}\preceq\chi I$ of (3.26) holds with $W=M(\hat{x},t)^{-1}$, $\bar{W}=\nu W$, $\nu=\overline{m}$, and $\chi=\overline{m}/\underline{m}$. As in Theorem LABEL:THM:Thm:CVSTEM:LMI, let $\beta$ be defined as $\beta=0$ for deterministic systems (4.3) and
@@
for stochastic systems (4.4), where $2\alpha_{e0}=L_{m}\bar{g}_{e0}^{2}(\alpha_{G}+{1}/{2})$, $2\alpha_{e1}=L_{m}\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}(\alpha_{G}+{1}/{2})$, $L_{m}$ is the Lipschitz constant of $\partial W/\partial x_{i}$, $\bar{g}_{e0}$ and $\bar{g}_{e1}$ are given in (4.4), and $\exists\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem LABEL:THM:Thm:robuststochastic.
@@
then Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic hold for the virtual systems (4.8) and (4.9), respectively, i.e., we have the following bounds for $\mathtt{e}=\hat{x}-x$ with $\nu=\overline{m}$ and $\chi=\overline{m}/\underline{m}$:
@@
where $V_{s\ell}=\int^{\hat{x}}_{x}\delta q^{\top}W\delta q$ and $V_{\ell}=\int^{\hat{x}}_{x}\|\Theta\delta q\|$ are given in Theorem LABEL:THM:Thm:path_integral with $W=M^{-1}=\Theta^{\top}\Theta$ defining a contraction metric, the disturbance bounds $\bar{d}_{e0}$, $\bar{d}_{e1}$, $\bar{g}_{e0}$, and $\bar{g}_{e1}$ are given in (4.3) and (4.4), respectively, $C_{e0}=\bar{g}_{e0}^{2}({2}{\alpha_{G}}^{-1}+1)$, and $C_{e1}=\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}({2}{\alpha_{G}}^{-1}+1)$. Note that for stochastic systems, the probability that $\|\mathtt{e}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Theorem LABEL:THM:Thm:CVSTEM:LMI indicates that (4.13) is equivalent to
@@
for $d$ in (4.8) and $G$ in (4.9), the bounds (4.14)–(4.16) follow from the proofs of Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic ncm, nscm. $\qed$
@@
Although (4.13) is not an LMI due to the nonlinear term $\nu\beta\mathrm{I}$ on its right-hand side for stochastic systems (4.4), it is a convex constraint as $\nu\beta=-\nu\alpha_{s}=-\nu\alpha_{e0}-\nu^{3}\alpha_{e1}$ is a concave function for $\nu\in\mathbb{R}_{>0}$ citeulike:163662, nscm.
4.2.2. CV-STEM Formulation for State Estimation
The estimator (4.7) gives a convex steady-state upper bound of the Euclidean distance between $x$ and $\hat{x}$ as in Theorem LABEL:THM:Thm:CVSTEMobjective mypaperTAC, ncm, nscm, mypaper.